# Gauge preservation in renormalization for Yang-Mills and gravity theories

### v2

@@ 2.5 The effective action and BRST symmetry
It is well known that the functional $G(J,\xi,\overline{\xi},K,L)$ allows to calculate the mean values
@@
The functional $\Gamma(<A>,<\eta>,<\overline{\eta}>,K,L)$ is known as the quantum effective action [49]-[51]. Of course, this action is also challenging to determine explicitly, as the mean values and $G(J,\xi,\overline{\xi},K,L)$ are typically difficult to be computed.
@@
$S_{gf}(A,\eta,\overline{\eta},K,L)=\int\;d^{4}x\;[L_{y}-\frac{\lambda}{2}F_{a}^{2}+(K_{\mu}^{a}-\overline{\eta}^{a}\phi^{ab}_{\mu})D^{\mu}_{ac}\eta^{c}+\frac{g}{2}L_{a}C_{abc}\eta^{b}\eta^{c}]$
@@
If the effective action $\Gamma(A,\eta,\overline{\eta},K,L)$ were determined, which is hard to achieve, then the functional $G_{\Gamma}(J,\xi,\overline{\xi},K,L)$ could be found perturbatively with the standard methods of Quantum Field Theory, by use of (2.50). The introduction of $\hbar$ explicitly implies the multiplication of every propagator by $\hbar^{-1}$ and every vertex by $\hbar$. Then, every graph is multiplied by $\hbar^{I-V}$ with $I$ the number of internal lines and $V$ the number of vertices. The quantity $I-V$ is related to the number of loops $L$ in a graph by the relation $I-V=L-1$. Therefore, an expansion of the form
$G_{\Gamma}(J,\xi,\overline{\xi},K,L)=\sum_{L=0}\hbar^{L-1}G^{(L)}_{\Gamma}(J,\xi,\overline{\xi},K,L)$
@@ 2.6 The Slavnov-Taylor equations and their implications
$L_{\text{counter}}=Z(\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu})[A_{\mu},A_{\nu}]$
with $Z$ a renormalization constant. This term alone does break gauge invariance and with gauge fixing terms it may break BRST invariance as well. At first sight, there may be no reason to care about BRST symmetry or gauge symmetry. After all, gauge invariance is not expected to hold if the gauge has been fixed and BRST may seem an accidental symmetry that plays no role. Instead, BRST symmetry is an important invariance. The action of the BRST on the ghosts on the other hand eliminate longitudinal models and enforce the Gauss law constraints [42]- [43] for the gauge theory. The same feature holds for gravity theories, as described for instance in section 2.2 of [44]. For these reaons BRST symmetry is something that one may desire to preserve order by order. Therefore, it is essential to understand the types of constraints that BRST invariance imposes on $\Gamma(A,\eta,\overline{\eta},K,L)$. This is the subject of the so-called Slavnov-Taylor identities [46]-[47].
@@
$e^{iG(J,\xi,\overline{\xi},K,L)}=\int DA_{\mu}^{a}D\eta^{a}D\overline{\eta}^{a}\exp\{i\int\;d^{4}x\;[L_{y}-\frac{\lambda}{2}F_{a}^{2}+(K_{\mu}^{a}-\overline{\eta}^{a}\phi^{ab}_{\mu})D^{\mu}_{ac}\eta^{c}-L_{a}s\eta^{a}$
@@
$\Gamma(A,\eta,\overline{\eta},K,L)=\Gamma(A,\eta,K_{\mu}^{a}-\overline{\eta}^{a}\phi^{ab}_{\mu},L),$ (2.58)
@@
$\Delta_{\mu}^{ab}=\alpha\partial_{\mu}\delta_{ab}-\beta gc_{abc}A^{b}_{\mu},\qquad d_{abc}=\gamma gc_{abc}.$
@@
The formula (2.59) is resemblant to (2.41). However, the difference is that the constants $\beta$ and $\gamma$ here are different in principle and that $\alpha$ is not necessarily equal to one. This difference appears since the BRST invariance and the consequent nilpotency conditions were not applied. As before $l(A_{\mu}^{a})$ has dimensions four or less and contains derivatives of the field, but it can not be assumed that is gauge invariant. It may contain terms of the form $\partial_{\mu}A^{a}_{\nu}F_{\mu\nu}^{a}$ or $A_{\mu}^{a}A_{\mu}^{a}$ and so on. In fact, its explicit expression will be found in the next sections.
@@ 2.7 Constraining the effective action by the Slavnov-Taylor identities
can be analyzed by use of the following formulas, which are derived from (2.59) and (2.60)
@@
$D^{\mu}_{(0)ab}\frac{\delta s_{h}(A_{0})}{\delta A^{\mu}_{0b}}=0,$
@@
$s_{p}=-c\int d^{4}xF_{\mu\nu}^{a}D_{\mu}A_{\nu}^{a}=-\frac{c}{2}\int d^{4}xF_{\mu\nu}^{a}(F^{a}_{\mu\nu}-gc_{abc}A^{b}_{\mu}A_{\nu}^{c})=2cS_{y}+\Delta S.$ (2.68)
In other words, this action corresponds to a lagrangian which is the sum of the Yang-Mills one $2cL_{y}(A)$ and $2\Delta L=cgc_{abc}F_{\mu\nu}^{a}A^{b}_{\mu}A_{\nu}^{c}$. The Euler-Lagrange expressions resulting from the extra piece are of first order and reads as
$\frac{\delta\Delta S}{\delta A_{\mu}^{a}}=2cgc_{abc}[(\partial_{\mu}A^{b}_{\nu})A_{\nu}^{c}-2(\partial_{\nu}A^{b}_{\mu})A_{\nu}^{c}+(\partial_{\nu}A^{b}_{\nu})A_{\mu}^{c}]-4cg^{2}c_{abc}c_{bde}A^{d}_{\mu}A^{e}_{\nu}A^{c}_{\nu}.$ (2.69)
@@
$D^{\mu}_{(0)ab}\frac{\delta S_{y}(A_{0})}{\delta A^{\mu}_{0b}}=0.$
@@
$D^{\mu}_{(0)ab}\frac{\delta\Delta S(A_{0})}{\delta A^{\mu}_{0b}}=2cgc_{abc}\partial_{\mu}[(\partial_{\mu}A^{b}_{0\nu})A_{0\nu}^{c}-2(\partial_{\nu}A^{b}_{0\mu})A_{0\nu}^{c}+(\partial_{\nu}A^{b}_{0\nu})A_{0\mu}^{c}]+O(g^{2})$
$=2gc_{abc}[(\partial_{\mu}A^{b}_{0\nu})\partial_{\mu}A_{0\nu}^{c}-2(\partial_{\nu}A^{b}_{0\mu})\partial_{\mu}A_{0\nu}^{c}+(\partial_{\nu}A^{b}_{0\nu}\partial_{\mu}A_{0\mu}^{c}+(\partial_{\mu}\partial_{\mu}A^{b}_{0\nu})\partial_{\mu}A_{0\nu}^{c}$
$-2(\partial_{\mu}\partial_{\nu}A^{b}_{0\mu})A_{\nu}^{c}+(\partial_{\mu}\partial_{\nu}A^{b}_{\nu})A_{0\mu}^{c}]+O(g^{2}).$
@@
$D^{\mu}_{(0)ab}\frac{\delta\Delta S(A_{0})}{\delta A^{\mu}_{0b}}=-2cgc_{abc}\partial_{\nu}[\partial_{\nu}A^{b}_{0\nu}-\partial_{\mu}A^{b}_{0\nu}]A_{0\mu}^{c}+O(g^{2}).$ (2.72)
@@
$g(\beta_{1}-\alpha_{1})c_{abc}A^{\mu}_{0c}\frac{\delta S_{y}(A_{0})}{\delta A^{\mu}_{0b}}=g(\beta_{1}-\alpha_{1})c_{abc}\partial_{\nu}[\partial_{\nu}A^{b}_{0\nu}-\partial_{\mu}A^{b}_{0\nu}]A_{\mu}^{c}+O(g^{2}).$ (2.73)
@@
$+g(\beta_{1}-\alpha_{1})(K-\overline{\eta}\phi)^{\mu}_{0a}c_{\mu ab}A_{0c\mu}\eta_{0b}+\frac{\beta_{1}g}{2}c_{abc}L_{0a}\eta_{0b}\eta_{0c}].$ (2.74)
@@ 2.8 Multiplicative renormalization
$+g(\beta_{1}-\alpha_{1})(K-\overline{\eta}\phi)^{\mu}_{0a}C_{\mu ab}A_{0c\mu}\eta_{0b}+\frac{\beta_{1}g}{2}C_{abc}L_{0a}\eta_{0b}\eta_{0c}].$ (2.78)
@@
$+S_{0gf}\bigg{(}(1+Z_{1}^{A})A_{0},(1+Z_{1}^{\eta})\eta_{0},Z_{1}^{\eta}\overline{\eta}_{0},(1+Z_{1}^{\eta})K_{0},(1+Z_{1}^{A})L_{0},(1+Z_{1}^{g})g_{0}\bigg{)}+O(\hbar^{2}).$ (2.81)
@@
$S_{1gf}=S_{0gf}\bigg{(}(1-Z_{1d}^{A})A_{0},(1-Z_{1d}^{\eta})\eta_{0},(1-Z_{1d}^{\eta})\overline{\eta}_{0},(1-Z_{1d}^{\eta})K_{0},(1-Z_{1d}^{A})L_{0},(1-Z_{1d}^{g})g_{0}\bigg{)}+O(\hbar^{2}),$ (2.89)
@@
$Z^{A}_{1}=1+\beta_{f}-\alpha_{f}+\frac{a_{f}}{2},\qquad Z_{1}^{\eta}=1+\frac{\alpha_{f}}{2},\qquad Z_{1}^{g}=1+\frac{a_{f}}{2}.$
@@
In brief, the discussion given above shows that the effective action is renormalized by multiplying the fields by certain constants. These multiplication constants are not all independent, some of them are identical to others as a consequence of the Slavnov-Taylor identities. All linear symmetries of the original action are maintained as symmetries of the effective actions. Even though the BRST symmetry is not exactly preserved, it remains valid order by order, up to higher-order terms in $\hbar$.
@@ 3.1 Diffeomorphisms in gravity theories
$ds=g_{\mu\mu}(y)dy^{\mu}dy^{\nu},$
in a curved variety $M_{4}$. The classical equations of motion for the model are assumed to be covariant under generic coordinate transformations. The most studied example of such type of theories is GR. Under general coordinate transformations $y^{\mu}\to x^{\mu}(y^{\nu})$, the metric tensor $g_{\mu\nu}$ is replaced by $\widetilde{g}_{\mu\nu}$
@@
The change under infinitesimal diffeomorphisms for $h^{\mu\nu}$ can be found by use of (3.96) by the following procedure. First note that
@@ 3.2 The non renormalizability of GR and renormalizable generalizations
Here $R(g_{\mu\nu})$ is the curvature tensor. This tensor may be expanded order by order in $h^{\mu\nu}$ by use of formulas such as (3.97)-(3.100) and higher order analogs. For instance
$\Gamma_{\rho\mu\nu}^{\alpha}=\frac{1}{2}\eta^{\alpha\beta}(\partial_{\mu}h_{\beta\nu}+\partial_{\nu}h_{\beta\mu}-\partial_{\beta}h_{\mu\nu})+O(h^{2}),$
@@
The expansion in $h^{\mu\nu}$ leads to non interacting terms, plus interacting terms from where the interaction vertices can be readed. The non interacting terms are collected in the so called zero order action, which has a generic form
@@
Here $D_{\mu\nu,\alpha\beta}$ is the standard propagator, and the last zeroth order action is written in terms of its inverse. The propagator contains second order derivatives in GR. This implies that the quantity $D^{-1}_{\mu\nu,\alpha\beta}$, in momentum space and for GR, goes as $k^{-2}$ after appropriate gauge fixing. In addition, the higher order terms give rise to $n$-graviton vertices, with $n$ being any possible integer. This is different for Yang-Mills theories, where there are only three and four vertex interactions for the gauge fields.

### v3

@@ 2.5 The effective action and BRST symmetry
It is well known that the functional $G(J,\xi,\overline{\xi},K,L)$ defined in (2.33) is a fundamental tool for the calculation of the mean values
@@
The functional $\Gamma(<A>,<\eta>,<\overline{\eta}>,K,L)$ is known as the quantum effective action [49]-[51]. Of course, the explicit determination of this action is a challenging task, as the mean values and $G(J,\xi,\overline{\xi},K,L)$ are typically difficult to be computed.
@@
$S_{gf}(A,\eta,\overline{\eta},K,L)=\int\;d^{4}x\;[L_{y}-\frac{\lambda}{2}F_{a}^{2}+(K_{\mu}^{a}-\overline{\eta}^{b}\phi^{ba}_{\mu})D^{\mu}_{ac}\eta^{c}+\frac{g}{2}L_{a}C_{abc}\eta^{b}\eta^{c}]$
@@
If it was possible to calculate exactly the effective action $\Gamma(A,\eta,\overline{\eta},K,L)$, which is a task hard to achieve, then the functional $G_{\Gamma}(J,\xi,\overline{\xi},K,L)$ could be found perturbatively with the standard methods of Quantum Field Theory, by use of (2.50). The introduction of $\hbar$ explicitly implies the multiplication of every propagator by $\hbar^{-1}$ and every vertex by $\hbar$. Then, every graph is multiplied by $\hbar^{I-V}$ with $I$ the number of internal lines and $V$ the number of vertices. The quantity $I-V$ is related to the number of loops $L$ in a graph by the relation $I-V=L-1$. Therefore, an expansion of the form
$G_{\Gamma}(J,\xi,\overline{\xi},K,L)=\sum_{L=0}\hbar^{L-1}G^{(L)}_{\Gamma}(J,\xi,\overline{\xi},K,L),$
@@ 2.6 The Slavnov-Taylor equations and their implications
$L_{\text{counter}}=Z(\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu})[A^{\mu},A^{\nu}],$
with $Z$ a renormalization constant. This term alone does break gauge invariance and with gauge fixing terms it may break BRST invariance as well.
At first sight, there may be no reason to care about BRST symmetry or gauge symmetry. After all, gauge invariance is not expected to hold if the gauge has been fixed and BRST may seem an accidental symmetry that plays no role. Instead, BRST symmetry is an important invariance. The action of the BRST on the ghosts on the other hand eliminate longitudinal models and enforce the Gauss law constraints [42]- [43] for the gauge theory. The same feature holds for gravity theories, as described for instance in section 2.2 of [44]. For these reasons BRST symmetry is something that one may desire to preserve order by order. Therefore, it is essential to understand the types of constraints that BRST invariance imposes on $\Gamma(A,\eta,\overline{\eta},K,L)$. This is the subject of the so-called Slavnov-Taylor identities [46]-[47].
@@
$e^{iG(J,\xi,\overline{\xi},K,L)}=\int DA_{\mu}^{a}D\eta^{a}D\overline{\eta}^{a}\exp\{i\int\;d^{4}x\;[L_{y}-\frac{\lambda}{2}F_{a}^{2}+(K_{\mu}^{a}-\overline{\eta}^{b}\phi^{ba}_{\mu})D^{\mu}_{ac}\eta^{c}-L_{a}s\eta^{a}$
@@
$\Gamma(A,\eta,\overline{\eta},K,L)=\Gamma(A,\eta,K_{\mu}^{a}-\overline{\eta}^{b}\phi^{ba}_{\mu},L),$ (2.58)
@@
$\Delta_{\mu}^{ab}=\alpha\partial_{\mu}\delta_{ab}-\beta gc_{abc}A^{c}_{\mu},\qquad d_{abc}=\gamma gc_{abc}.$
@@
The formula (2.59) is pretty close to (2.41). However, the difference is that the constants $\beta$ and $\gamma$ here are different in principle and that $\alpha$ is not necessarily equal to one. This difference appears since the BRST invariance and the consequent nilpotency conditions were not applied. As before $l(A_{\mu}^{a})$ has dimensions four or less and contains derivatives of the field, but it can not be assumed that is gauge invariant. It may contain terms of the form $\partial_{\mu}A^{a}_{\nu}F^{a\mu\nu}$ or $A_{\mu}^{a}A^{\mu}_{a}$ and so on. In fact, its explicit expression will be found in the next sections.
@@ 2.7 Constraining the effective action by the Slavnov-Taylor identities
can be analyzed by use of the following formulas, which are derived diretly from (2.59) and (2.60)
@@
$D^{\mu}_{(0)ab}\frac{\delta s_{h}(A_{0})}{\delta A^{\mu}_{0b}}=0.$
@@
$s_{p}=-c\int d^{4}xF^{a\mu\nu}D_{\mu}A_{\nu}^{a}=-\frac{c}{2}\int d^{4}xF^{a\mu\nu}(F^{a}_{\mu\nu}-gc_{abc}A^{b}_{\mu}A_{\nu}^{c})=2cS_{y}+\Delta S.$ (2.68)
In other words, this action corresponds to a lagrangian which is the sum of the Yang-Mills one $2cL_{y}(A)$ and $2\Delta L=cgc_{abc}F_{\mu\nu}^{a}A^{b\mu}A^{c\nu}$. The Euler-Lagrange expressions resulting from the extra piece are of first order and reads as
$\frac{\delta\Delta S}{\delta A_{\mu}^{a}}=2cgc_{abc}[(\partial_{\mu}A^{b}_{\nu})A^{c\nu}-2(\partial_{\nu}A^{b}_{\mu})A^{c\nu}+(\partial_{\nu}A^{b\nu})A_{\mu}^{c}]-4cg^{2}c_{abc}c_{bde}A^{d}_{\mu}A^{e}_{\nu}A^{c\nu}.$ (2.69)
@@
$D^{\mu}_{(0)ab}\frac{\delta S_{y}(A_{0})}{\delta A^{\mu}_{0b}}=0,$
@@
$D^{\mu}_{(0)ab}\frac{\delta\Delta S(A_{0})}{\delta A^{\mu}_{0b}}=2cgc_{abc}\partial^{\mu}[(\partial_{\mu}A^{b}_{0\nu})A_{0}^{c\nu}-2(\partial_{\nu}A^{b}_{0\mu})A_{0}^{c\nu}+(\partial_{\nu}A^{b\nu}_{0})A_{0\mu}^{c}]+O(g^{2})$
$=2gc_{abc}[(\partial_{\mu}A^{b}_{0\nu})(\partial^{\mu}A_{0}^{c\nu})-2(\partial_{\nu}A^{b}_{0\mu})(\partial^{\mu}A_{0}^{c\nu})+(\partial_{\nu}A^{b\nu}_{0})(\partial_{\mu}A_{0}^{c\mu})+(\partial^{\mu}\partial_{\mu}A^{b}_{0\nu})A_{0}^{c\nu}$
$-2(\partial^{\mu}\partial_{\nu}A^{b}_{0\mu})A_{0}^{c\nu}+(\partial^{\mu}\partial_{\nu}A^{b\nu}_{0})A_{0\mu}^{c}]+O(g^{2}).$
@@
$D^{\mu}_{(0)ab}\frac{\delta\Delta S(A_{0})}{\delta A^{\mu}_{0b}}=-2cgc_{abc}\partial^{\nu}[\partial_{\mu}A^{b}_{0\nu}-\partial_{\nu}A^{b}_{0\mu}]A_{0}^{c\mu}+O(g^{2}).$ (2.72)
@@
$g(\beta_{1}-\alpha_{1})c_{abc}A^{\mu}_{0c}\frac{\delta S_{y}(A_{0})}{\delta A^{\mu}_{0b}}=g(\beta_{1}-\alpha_{1})c_{abc}\partial^{\nu}[\partial_{\mu}A^{b}_{0\nu}-\partial_{\nu}A^{b}_{0\mu}]A_{0}^{c\mu}+O(g^{2}).$ (2.73)
@@
$+g(\beta_{1}-\alpha_{1})(K-\overline{\eta}\phi)^{\mu}_{0a}c_{abc}A_{0c\mu}\eta_{0b}+\frac{\beta_{1}g}{2}c_{abc}L_{0a}\eta_{0b}\eta_{0c}].$ (2.74)
@@ 2.8 Multiplicative renormalization
$+g(\beta_{1}-\alpha_{1})(K-\overline{\eta}\phi)^{\mu}_{0a}C_{abc}A_{0c\mu}\eta_{0b}+\frac{\beta_{1}g}{2}C_{abc}L_{0a}\eta_{0b}\eta_{0c}].$ (2.78)
@@
$+S_{0gf}\bigg{(}(1+Z_{1}^{A})A_{0},(1+Z_{1}^{\eta})\eta_{0},Z_{1}^{\eta}\overline{\eta}_{0},(1+Z_{1}^{\eta})K_{0},(1+Z_{1}^{A})L_{0},(1+Z_{1}^{g})g_{0}\bigg{)}+O(\hbar^{2}),$ (2.81)
@@
$S_{1gf}=S_{0gf}\bigg{(}(1-Z_{1d}^{A})A_{0},(1-Z_{1d}^{\eta})\eta_{0},(1-Z_{1d}^{\eta})\overline{\eta}_{0},(1-Z_{1d}^{\eta})K_{0},(1-Z_{1d}^{A})L_{0},(1-Z_{1d}^{g})g_{0}\bigg{)}+O(\hbar^{2}).$ (2.89)
@@
$Z^{A}_{1}=1+\beta_{f}-\alpha_{f}+\frac{a_{f}}{2},\qquad Z_{1}^{\eta}=1+\frac{\alpha_{f}}{2},\qquad Z_{1}^{g}=1+\frac{a_{f}}{2}.$ (2.94)
@@
In brief, the discussion given above shows that the effective action is renormalized by multiplying the fields by certain constants. These multiplication constants are not all independent (2.94), some of them are identical to others as a consequence of the Slavnov-Taylor identities. All linear symmetries of the original action are maintained as symmetries of the effective actions. Even though the BRST symmetry is not exactly preserved, it remains valid order by order, up to higher-order terms in $\hbar$.
@@ 3.1 Diffeomorphisms in gravity theories
$ds^{2}=g_{\mu\mu}(y)dy^{\mu}dy^{\nu},$
in a curved space $M_{4}$. The classical equations of motion for the model are assumed to be covariant under generic coordinate transformations. The most studied example of such type of theories is GR. Under general coordinate transformations $y^{\mu}\to x^{\mu}(y^{\nu})$, the metric tensor $g_{\mu\nu}$ is replaced by $\widetilde{g}_{\mu\nu}$
@@
The change under infinitesimal diffeomorphisms for $h^{\mu\nu}$ can be found by use of (3.97) by the following procedure. First employ the formula
@@ 3.2 The non renormalizability of GR and renormalizable generalizations
Here $R(g_{\mu\nu})$ is the curvature scalar. Recall that the Riemann curvature tensor, in a local system of coordinates, can be calculated in terms of the Christofel symbols
$\Gamma_{\mu\nu}^{\alpha}=\frac{1}{2}g^{\alpha\beta}(\partial_{\mu}g_{\beta\nu}+\partial_{\nu}g_{\beta\mu}-\partial_{\beta}g_{\mu\nu}),$
as follows
$R^{\beta}_{\alpha\mu\nu}=\partial_{\mu}\Gamma_{\nu\alpha}^{\beta}-\partial_{\nu}\Gamma_{\mu\alpha}^{\beta}+\Gamma_{\mu\gamma}^{\beta}\Gamma_{\nu\alpha}^{\gamma}-\Gamma_{\nu\gamma}^{\beta}\Gamma_{\mu\alpha}^{\gamma}.$
This tensor may be expanded order by order in $h^{\mu\nu}$ by use of formulas such as (3.98)-(3.101) and higher order analogs. For instance
$\Gamma_{\mu\nu}^{\alpha}=\frac{1}{2}\eta^{\alpha\beta}(\partial_{\mu}h_{\beta\nu}+\partial_{\nu}h_{\beta\mu}-\partial_{\beta}h_{\mu\nu})+O(h^{2}),$
@@
The expansion in $h^{\mu\nu}$ leads to non interacting terms, plus interacting ones from where the interaction vertices can be read. The non interacting terms are collected in the so called zero order action, which has a generic form
@@
Here $D_{\mu\nu,\alpha\beta}$ is the standard propagator, and the last zeroth order action is written in terms of its inverse. The propagator in standard GR contains second order derivatives. This implies that the quantity $D^{-1}_{\mu\nu,\alpha\beta}$, in momentum space and for GR, goes as $k^{-2}$ after appropriate gauge fixing. In addition, the higher order terms give rise to $n$-graviton vertices, with $n$ being any possible integer. This is different for Yang-Mills theories, where there are only three and four vertex interactions for the gauge fields.